OpenStudy (kimberly_pr):

Which of the following is sufficient to establish that wxyz = 0?

OpenStudy (kimberly_pr):

w=y-z xz=0 xy-yz=0 wx=zy

jimthompson5910 (jim_thompson5910):

hint: wxyz = 0 is the same as (wy)*(xz) = 0 after you rearrange a few terms

OpenStudy (kimberly_pr):

So b? Or d?... I'm still a little confused

jimthompson5910 (jim_thompson5910):

if (wy)*(xz) = 0, then either factor is zero (or both could be 0)

jimthompson5910 (jim_thompson5910):

if A*B = 0, then either A = 0 or B = 0 (or both are zero at the same time)

OpenStudy (kimberly_pr):

SO I guess B?

jimthompson5910 (jim_thompson5910):

yes wy = 0 or xz=0

OpenStudy (zzr0ck3r):

0 times anything is 0